OpenStudy (anonymous):

Is 3sec(x) 3/sin(x) or 1/3sin(x) ?

OpenStudy (anonymous):

3sec(x) = 3/cos(x)

OpenStudy (kc_kennylau):

\[3\sec x=3\times\frac1{\cos x}=\frac3{\cos x}\]

OpenStudy (anonymous):

Ah, THANK YOU so much everyone :)

OpenStudy (kc_kennylau):

no problem :)